OpenStudy (2yoututu11):

help A pastry recipe calls for 3 cups of flour to make 8 servings. How many cups of flour are needed to make 6 servings?

OpenStudy (s):

3/8 = x/6 3*6 = x*8 18 = 8x x = 18/8 x = 2.25 cups are needed to make 6 servings